# Beauty and the Bias: Exploring the Impact of Attractiveness on Multimodal Large Language Models

### v2

@@ Appendix F Gender Bias in the Gender Stereotyped Jobs Scenarios
Tables 17 - 23 below detail the strength of the gender bias in the gender stereotyped jobs across all tested models. Each table reports the test statistic from the Kruskal–Wallis test (${H}_{i}^{\textit{gender}}$) comparing the values of $\phi_{i}$ between male and female images for each scenario $s_{i}$. Standard star notation is employed to indicate the level of statistical significance: ***∗ ∗ ∗ denotes $p<0.001$ and **∗ ∗ denotes $p<0.01$. Additionally, the tables provide the mean values of $\phi_{i}$ for both male and female images. Scenarios exhibiting statistically significant differences are highlighted in bold, with emphasis on the group displaying the higher mean value, indicating a stronger model tendency to associate that group with the “Stereotyped Choice” (i.e., Choice 1). This pattern reflects the model’s underlying preference or bias. Notably, across nearly all scenarios, the model demonstrates a consistent inclination to associate images of males with jobs traditionally associated with males thereby replicating existing societal gender stereotypes, even though gender should not inform the likelihood of the person performing either one of these jobs.
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x)$: male Mean $\text{Mean }\phi_{i}(x)$: female ${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x)$: male Mean $\text{Mean }\phi_{i}(x)$: female ${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x)$: male Mean $\text{Mean }\phi_{i}(x)$: female ${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x)$: male Mean $\text{Mean }\phi_{i}(x)$: female ${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x)$: male Mean $\text{Mean }\phi_{i}(x)$: female ${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x)$: male Mean $\text{Mean }\phi_{i}(x)$: female ${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x)$: male Mean $\text{Mean }\phi_{i}(x)$: female ${H}_{i}^{\textit{gender}}$
@@ Appendix G Racial Bias in the Race Stereotyped Jobs Scenarios
Standard star notation is employed to indicate the level of statistical significance: ***∗ ∗ ∗ denotes $p<0.001$ and **∗ ∗ denotes $p<0.01$. Scenarios exhibiting statistically significant differences are highlighted in bold, with emphasis on the group displaying the higher mean value, indicating a stronger model tendency to associate that group with the first choice. This pattern reveals the model’s underlying preferences or biases. Notably, instances of race bias are observed across multiple scenarios. In nearly all such cases, the direction of the bias aligns with prevailing societal racial stereotypes, suggesting that the model not only internalizes but also reproduces these stereotypes in its decision-making processes.
Choice 1 Choice 2 Mean $\textit{Mean }\phi_{i}(x)$: Race 1 Mean $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 Mean $\textit{Mean }\phi_{i}(x)$: Race 1 Mean $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 Mean $\textit{Mean }\phi_{i}(x)$: Race 1 Mean $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 Mean $\textit{Mean }\phi_{i}(x)$: Race 1 Mean $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 Mean $\textit{Mean }\phi_{i}(x)$: Race 1 Mean $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 Mean $\textit{Mean }\phi_{i}(x)$: Race 1 Mean $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 Mean $\textit{Mean }\phi_{i}(x)$: Race 1 Mean $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$

### v3

@@ Appendix F Gender Bias in the Gender Stereotyped Jobs Scenarios
Tables 17 - 23 below detail the strength of the gender bias in the gender stereotyped jobs across all tested models. Each table reports the test statistic from the Kruskal–Wallis test (${H}_{i}^{\textit{gender}}$) comparing the values of $\phi_{i}$ between male and female images for each scenario $s_{i}$. Standard star notation is employed to indicate the level of statistical significance: *** denotes $p<0.001$ and ** denotes $p<0.01$. Additionally, the tables provide the mean values of $\phi_{i}$ for both male and female images. Scenarios exhibiting statistically significant differences are highlighted in bold, with emphasis on the group displaying the higher mean value, indicating a stronger model tendency to associate that group with the “Stereotyped Choice” (i.e., Choice 1). This pattern reflects the model’s underlying preference or bias. Notably, across nearly all scenarios, the model demonstrates a consistent inclination to associate images of males with jobs traditionally associated with males thereby replicating existing societal gender stereotypes, even though gender should not inform the likelihood of the person performing either one of these jobs.
Choice 1 Choice 2
$\text{Mean }\phi_{i}(x)$
$\text{Mean }\phi_{i}(x)$
${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2
$\text{Mean }\phi_{i}(x)$
$\text{Mean }\phi_{i}(x)$
${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2
$\text{Mean }\phi_{i}(x)$
$\text{Mean }\phi_{i}(x)$
${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2
$\text{Mean }\phi_{i}(x)$
$\text{Mean }\phi_{i}(x)$
${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2
$\text{Mean }\phi_{i}(x)$
$\text{Mean }\phi_{i}(x)$
${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2
$\text{Mean }\phi_{i}(x)$
$\text{Mean }\phi_{i}(x)$
${H}_{i}^{\textit{gender}}$
@@
Choice 1 Choice 2
$\text{Mean }\phi_{i}(x)$
$\text{Mean }\phi_{i}(x)$
${H}_{i}^{\textit{gender}}$
@@ Appendix G Racial Bias in the Race Stereotyped Jobs Scenarios
Standard star notation is employed to indicate the level of statistical significance: *** denotes $p<0.001$ and ** denotes $p<0.01$. Scenarios exhibiting statistically significant differences are highlighted in bold, with emphasis on the group displaying the higher mean value, indicating a stronger model tendency to associate that group with the first choice. This pattern reveals the model’s underlying preferences or biases. Notably, instances of race bias are observed across multiple scenarios. In nearly all such cases, the direction of the bias aligns with prevailing societal racial stereotypes, suggesting that the model not only internalizes but also reproduces these stereotypes in its decision-making processes.
Choice 1 Choice 2 $\textit{Mean }\phi_{i}(x)$: Race 1 $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 $\textit{Mean }\phi_{i}(x)$: Race 1 $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 $\textit{Mean }\phi_{i}(x)$: Race 1 $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 $\textit{Mean }\phi_{i}(x)$: Race 1 $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 $\textit{Mean }\phi_{i}(x)$: Race 1 $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 $\textit{Mean }\phi_{i}(x)$: Race 1 $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$
@@
Choice 1 Choice 2 $\textit{Mean }\phi_{i}(x)$: Race 1 $\textit{Mean }\phi_{i}(x)$: Race 2 ${H}_{i}^{\textit{race}}$